# Practical privacy metrics for synthetic data

### v4

@@ 5 Excluding records
Three other variables are flagged as having disclosive two_way relationships workclass, marital, relationship with totals of 7,7 and 2 $tq$ combinations respectively. The summary function multi.disclosure does not allow target and key specific pairs to be excluded. To investigate this it is necessary to examine the output of disclosure151515see code in Appendix 3 for how to do this.. This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@ 7 Acknowledgement
Research Data Scotland (https://www.researchdata.scot/) funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package161616from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control. We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14. We would also like to thank Timon Huijser, for pointing out errros in a previous version of this vignette.
Appendix 1: Notation and formal definitions
@@
For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ is the count of records with this $q$ and level t=1,Tt=1,...T of the target. The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$. When a member of $q$ is in the original data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero. Similarly when a member of $q$ is in the SD but not in the original, $d_{.q}$ and $d_{tq}$ are all zero. The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~only}$ and $N_{d~both}$. The equivalent totals for the SD are $N_{s}$, $N_{s~only}$ and $N_{s~both}$.
@@
The measures $DCAP$ and $DCAP_{alt}$ are each based on the percentage of synthetic records where $t$ is correctly predicted in the SD This gives
$DCAP=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{sboth}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{s}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{d}$
The $TCAP$ and $TCAP_{alt}$ measure use the same count of disclosive records as $DiSCO$ in their numerators, but use different denominators. $TCAP_{alt}$ was mistakenly quoted in earlier versions of this vignette as being the measure proposed in [11], using $N_{s~both}$ as its denominator:
$TCAP_{alt}=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{s~both}.$ (10)
but the measure used in [11, 10] had a smaller denominator restricted to all counts in the original datam with keys corresponding to those predicted with certainty from the SD, for any target:
@@
where the summation over $t$ in the denominator indicates columns of the $tq$ tables where any target is predicted with certainty from the SD,
See section 3.3.2 for examples of these measures.

### v5

@@ 5 Excluding records
Three other variables are flagged as having disclosive two_way relationships workclass, marital, relationship with totals of 7,7 and 2 $tq$ combinations respectively. The summary function multi.disclosure does not allow target and key specific pairs to be excluded. To investigate this it is necessary to examine the output of disclosure. This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@ 7 Acknowledgement
Research Data Scotland (https://www.researchdata.scot/) funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package131313from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control. We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14. We would also like to thank Timon Huijser, for pointing out errros in a previous version of this vignette.
Appendix
Notation and formal definitions
@@
For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ is the count of records with this $q$ and level t=1,Tt=1,...T of the target. The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$. When a member of $q$ is in the original data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero. Similarly when a member of $q$ is in the SD but not in the original, $d_{.q}$ and $d_{tq}$ are all zero. The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~only}$ and $N_{d~both}$. The equivalent totals for the SD are $N_{s}$, $N_{s~only}$ and $N_{b}$.
@@
The measures $DCAP_{b}$ and $DCAP_{s}$ are each based on the percentage of synthetic records where $t$ is correctly predicted in the SD This gives
$DCAP_{b}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{b}~~~DCAP_{s}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{s}$
Another option for the denominator of a $DCAP$ measure is the count of GT records, $N_{d}$, as used for $DiSCO$ in equation (11).
$DCAP_{d}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{d}$ (10)
When the SD and GT have the same number of records $DCAP_{s}=DCAP_{d}$. But when the SD has fewer records $DCAP_{d}$ will be reduced.
The $TCAP$ and $TCAP_{b}$ measures use the same count of disclosive records as $DiSCO$ in their numerators, but use different denominators. $TCAP_{b}$ was mistakenly quoted in earlier versions of this vignette as being the measure proposed in [11], using $N_{b}$ as its denominator:
$TCAP_{b}=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{b}.$ (11)
but the measure used in [11, 10] had a smaller denominator restricted to all counts in the original data, with keys corresponding to those predicted with certainty from the SD, for any target:
@@
where the summation over $t$ in the denominator indicates columns of the $tq$ tables where any target is predicted with certainty from the SD, the quantity $N_{bp}$ defined in section 3.3.2.